The horizontal velocity of Bruce’s Fastpitch is 40 m/s at the instant release from his hand if the horizontal distance from Bruce’s hand to the Homeplate is 17.5 m at the instance of release how much time does a batter have to react to the pitch and swing the bat
Answer #1

Given

s = 17.5 m

v = 40 m/s

As we know ,

t = s/v = 17.5 / 40 = 0.4375 second

So batter is having a time of 0.4375 second